OpenStudy (taepoohwae):

Can someone please show the steps for this and give the ordered pair? use the method of substitution to solve the system of linear equations. 2x-y=4 7x-3y=16

OpenStudy (anonymous):

Solve for y or x Plug in that value of y or x into the other equation (in parenthesis) The other equation has only 1 variable now, solve for it. Put that result into the first equation to get y or x.

OpenStudy (taepoohwae):

I'm new to all of this. Therefore I don't understand, and need to be shown step by step to solve completely.

OpenStudy (campbell_st):

take the 1st equation and make y the subject... can you do that..?

OpenStudy (taepoohwae):

use the method of substitution to solve the system of linear equations. 2x-y=4 7x-3y=16

OpenStudy (campbell_st):

so do you know how to make y the subject in the equation 2x - y = 4

OpenStudy (taepoohwae):

As I mentioned this is my first algebra class so I know nothing. I need someone to solve this step by step.

OpenStudy (campbell_st):

ok... so add y to both sides of the equation 2x = 4 + y now subtract 4 from both sides 2x - 4 = y now this can be substituted into the 2nd equation inplace of y 7x - 3(2x - 4) = 16 distributing 7x - 6x + 12 = 16 solve for x when you have x substitute it into the 1st equation and solve for y

OpenStudy (taepoohwae):

Thanks still not sure I really understand, but I'll study this more and use it as my notes.